5. Consider the following basket of student essentials:

2019

2020

Burritos

$6.30

$6.60

Pizzas

$9.90

$10.00

Movie Tickets

$10.00

$10.50

Energy Drinks

$4.99

$5.99

a. Using 2019 as the base year, what is the CPI in 2019?

b. Assume students buy 30 burritos, 25 pizzas, 15 movie tickets, and 200 energy drinks each year. What is the CPI in 2020 using 2019 as a base year? What is the inflation rate? Would this rate be higher or lower if students bought more energy drinks?
Answer #1
As per the information given in the question
Goods Year 2019
Price (P) in $

Quantity (Q) in 2019
Expenditure in 2019 in $ = P x Q Year 2020
Price (P) in $

Quantity (Q) in 2020
Expenditure in 2020 in $ = P x Q
Burritos 6.30 30 189 6.60 30 198
Pizzas 9.90 25 247.50 10.00 25 250
Movie Tickets 10.00 15 150 10.50 15 157.50
Energy Drinks 4.99 200 998 5.99 200 1198
1584.50 1803.50
Cost of Basket in 2019 = ∑PxQ = $1584.50
Cost of Basket in 2020 = ∑PxQ = $1803.50
(A) as per the question 2019 is the base year so the CPI of 2019 =100%
As because the CPI of the base year is always 100%
(B) CPI of 2020 = (Cost of Basket in 2020/Cost of Basket in 2019)x100 = (1803.50/1584.50)x100 = 113.82%
(C) Inflation Rate = {(CPI of 2020 - CPI of 2019)/CPI of 2019}x100
Inflation Rate = {(113.82 - 100)/100}x100= 13.82%
(D) As per the question if the students buy more energy drinks then the inflation rate will be higher
suppose instead of 200 now the students demand 250 of energy drinks then
Goods Year 2019
Price (P) in $

Quantity (Q) in 2019
Expenditure in 2019 in $ = P x Q Year 2020
Price (P) in $

Quantity (Q) in 2020
Expenditure in 2020 in $ = P x Q
Burritos 6.30 30 189 6.60 30 198
Pizzas 9.90 25 247.50 10.00 25 250
Movie Tickets 10.00 15 150 10.50 15 157.50
Energy Drinks 4.99 250 1247.50 5.99 250 1497.50
1834 2103
Cost of Basket in 2019 = ∑PxQ = $1834
Cost of Basket in 2020 = ∑PxQ = $2103
as per the question 2019 is the base year so the CPI of 2019 =100%
As because the CPI of the base year is always 100%
CPI of 2020 = (Cost of Basket in 2020/Cost of Basket in 2019)x100 = (2103/1834)x100 = 114.67%
Inflation Rate = {(CPI of 2020 - CPI of 2019)/CPI of 2019}x100
Inflation Rate = {(114.67 - 100)/100}x100= 14.67%
inflation rate calculated after buying more energy drinks is higher than the earlier inflation rate
so if the students buy more energy drinks then the inflation rate will be higher